OpenStudy (anonymous):

what is the solution to the system of equations x+y=2 x-y=4 when x=3 and y=-1

OpenStudy (anonymous):

We can use subsitution to solve this, our two equations are \[ x+y=2\]\[ x-y=4\] we can subtract x from the first equation to find what y equals.\[ x-x+y=2-x\]\[ y=2-x\] Now that we know what y is equal to we can subsitute 2-x for y. \[ x-y=4\]\[ x-(2-x)=4\]\[ x-2+x=4\]\[ 2x-2=4\]\[ 2x=6\]\[ x=3\]So we have for our solution so far (3 , y ). However we are not done yet. You can finish this on your own by same thing but instead plug in 3 for x.

OpenStudy (anonymous):

i still don't understand

OpenStudy (anonymous):

an easier method: elimination method